gianemarie
gianemarie gianemarie
  • 03-03-2020
  • Mathematics
contestada

Solve using normalcdf

Solve using normalcdf class=

Respuesta :

MathPhys
MathPhys MathPhys
  • 04-03-2020

Answer:

0.209

Step-by-step explanation:

Find the sample mean and standard deviation.

μ = 1050

s = 218 / √50 = 30.8

Find the z-score.

z = (1075 − 1050) / 30.8

z = 0.81

Find the probability.

P(Z > 0.81) = 1 − 0.7910 = 0.2090

Answer Link

Otras preguntas

What is the slope of any line parallel to the line 8x + 9y = 3 in the standard (x,y) coordinate plane?
Suppose you ride your bicycle to the library traveling at 0.5 km/min. It takes you 25 minutes to get to 1.3 the library. How far did you travel?
USE A NUMBER LINE TO FIND 8/9 / 8/45.
3x+y=4 2x+y=5 solve using linear combination
China, Korea, and Japan share cultural similarities in part due to their (1) clashes with Russian imperialists (2) shared river systems (3) contacts through tra
what does etymology study?
3x+y=4 2x+y=5 solve using linear combination
The significance of the Open Door Policy for the United States lay in its (A) providing a use for the Philippines as a way station between China and the United
What is 1997 from 2014????????????
What are abiotic factors of the enviroment lynx